OpenStudy (anonymous):

If LCM of two 6-digit integers has 10 digits, then their GCD has at most how many digits?

OpenStudy (rational):

\[\text{lcm(a,b)} \times \gcd(a,b) = a\times b\]

OpenStudy (rational):

The product of any two \(6\)-digit numbers can have\(11, 12, 13\) digits. Since the \(\text{lcm}\) has \(10\) digits, \(\gcd\) can have \(1, 2,3,4\) digits.

OpenStudy (anonymous):

how did you arrived at this conclusion that lcm has 10 digits so GCD at most has 4 digits

OpenStudy (rational):

Is it possible for the product of a 10 digit number and a 5 digit number to have 13 digits ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

but the answer 4 is not correct

OpenStudy (anonymous):

...as per to my laptop.

OpenStudy (mathmath333):

should be \(2\) then

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \text{lcm(a,b)} \times \gcd(a,b) = a\times b\hspace{.33em}\\~\\ \text{10 digit} \times \gcd(a,b) = \text{12 digit}\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

the product of two \(6\) digit natural numbes has \(12 ~~digits\)

OpenStudy (mathmath333):

http://www.wolframalpha.com/input/?i=999999*999999+%3D

OpenStudy (anonymous):

Ok. I see on website where you re coming from .. so max can have 12 digit on multiplying two 6 digits numbers.... but how did you arrived that if it were not for this website wolf...

OpenStudy (mathmath333):

u calculate the product largest 6 digit numbers by hand also

OpenStudy (anonymous):

hahha..... thought there was an easier way

OpenStudy (mathmath333):

but this doenst prove anything

OpenStudy (anonymous):

both ans. 2, 4, are wrong

OpenStudy (mathmath333):

how did u know that

OpenStudy (mathmath333):

only 1 option remains now

OpenStudy (anonymous):

Ans is 3 : In this case we know \[ $a<10^6$ and $b<10^6$, so\\ $ab<10^{12}$.\] We also know that \[$\mathop{\text{lcm}}[a,b]\ge 10^9$\] since the smallest 10-digit number is $10^9$ Therefore $$\gcd(a,b) < \frac{10^{12}}{10^9} = 10^3,$$ so $\gcd(a,b)$ has at most $\boxed{3}$ digits.

OpenStudy (anonymous):

\[so $\gcd(a,b)$ \\has at most\\ $\boxed{3}$ digits.\]

OpenStudy (mathmath333):

Ans is 3 : In this case we know \(a<10^6\) and \(b<10^6\), so \(ab<10^{12}\) We also know that \(lcm[a,b]≥10^9\) since the smallest 10-digit number is \(10^9\) Therefore \(\gcd(a,b)<\dfrac{10^{12}}{10^9}=10^3\) so \(\gcd(a,b)\) has at most\(\boxed{3}\) digits.

OpenStudy (mathmath333):

looks good now